# A case study of gas impacted by black-hole jets with the JWST: outflows, bow shocks, and high excitation of the gas in the galaxy IC5063

### v1

@@ Abstract
We present James Webb Space Telescope MIRI data of the inner 32kpc2 of the galaxy IC5063, in which the jets of a supermassive black hole interact with the gaseous disk they are crossing. Jet-driven outflows were known to be initiated along or near the jet path, and the stability conditions of clouds were known to vary because of these outflows. The MIRI data, of unprecedented resolution and sensitivity in the infrared, now reveal that there are more than ten discrete regions with outflows, nearly doubling the number of such known regions. Outflows exist near the radio lobes, at the nucleus, in a biconical structure perpendicular to the jet, and in a bubble moving against the disk. In some of them, velocities above escape velocity are observed. Stratification is also observed, with higher ionization or excitation gas attaining higher velocities. More outflows and bow shocks, found further away from the nucleus than the radio lobes, in regions without significant radio emission, reveal the existence of past or weak radio jets that interacted with the interstellar medium. The coincidence of the bow shocks with the optical extended emission line region (EELR) suggests that the jets also contributed to the gas ionization. Maps of the gas excitation temperature, Tex, indicate that the molecular gas is most excited in regions with radio emission. There, Tex is 100 K higher than in the EELR interior. We argue that a combination of jet-related shocks and cosmic rays is likely responsible for this excess molecular gas excitation.
@@ 1 Introduction
The advent of the James Webb Space Telescope (JWST), with its unprecedented spatial resolution and sensitivity in the infrared, enabled the astronomical community to make fast advancements in the field of galaxy evolution, by looking further in time for distant galaxies and deeper into nearby galaxies. Besides the detection of the first galaxies and even the first stars in them (e.g., Atek et al. 2022; Castellano et al. 2022; Naidu et al. 2022; Finkelstein et al. 2022, 2023; Donnan et al. 2023a, b; Harikane et al. 2023a; Leung et al. 2023; Maiolino et al. 2023a; Yan et al. 2023), a major galaxy evolution theme to be revised by the JWST is the impact of black holes (BHs) on galaxy evolution (including via outflows; Armus et al. 2022; Rupke et al. 2023; Vayner et al. 2024). Cycle 1 surveys already indicated that many more active galactic nuclei (AGN) exist early, above redshift 3-4 (Barro et al., 2023; Juodzbalis et al., 2023; Kocevski et al., 2023; Labbe et al., 2023; Lyu et al., 2023), compared to what was known from past observations in X-rays or other wavelengths, and that some black holes grew too fast too soon compared to the local BH mass vs. host galaxy relation (e.g., Harikane et al. 2023b; Larson et al. 2023; Maiolino et al. 2023b for 4z8.5).
@@
The turmoil that jets cause as they propagate through a dense ISM is demonstrated in this paper, with the aid of cycle 1 JWST data of the nearby, early-type galaxy IC5063 (taken for the program 2004; PI Dasyra). IC5063 is well suited for demonstrating the capabilities of the JWST, as it is a galaxy known for its numerous jet-cloud interactions. These interactions occur as radio jets cross a disk of dense gas for hundreds of parsecs, along a trajectory nearly parallel to the disk plane, prior to escaping it (Morganti et al., 1998). The jets collide with dense ISM clouds at the two radio lobes, 300-500 pc away from the nucleus, driving outflows. An HI outflow has been seen in absorption in front of one of the radio lobes (Oosterloo et al., 2000; Morganti et al., 2007), and in emission in several regions, in tracers of all gas phases - atomic (Morganti et al., 2007; Dasyra et al., 2015; Congiu et al., 2017; Venturi et al., 2021) and molecular (Tadhunter et al., 2014; Morganti et al., 2015; Dasyra et al., 2015, 2016, 2022). In Very Large Telescope (VLT) SINFONI data, Dasyra et al. (2015) found numerous outflows starting at or near the jet trail: at least 6 discrete outflow starting points were reported for the central 22kpc2 of this galaxy, including at the two radio lobes. In some of these locations, gas moving faster than the escape velocity was detected. Moreover, the gas in nearly all of the inner 1.5 square kiloparsec of the galaxy shows distorted kinematic signatures due to the jet. In VLT MUSE data, Venturi et al. (2021) showed that, in the ionized gas as seen by [O III], the turbulence in the galaxy may reach 4 kpc in both directions perpendicular to the jet. Morganti et al. (2015) proposed even the propagation of a homonegeous jet-inflated cavity (a.k.a. cocoon) in the inner region of IC5063 based on the low rotational number J CO data distribution. The mass of all outflowing gas is at least 6106 , when taking into account the mass of the neutral gas and the cold molecular gas, using their most-abundant gas tracers, HI and CO (Morganti et al., 2007; Dasyra et al., 2016). The mass of the outflowing cold required a CO luminosity to mass factor different than Galactic (Dasyra et al. 2016; see also Bolatto et al. 2013), because dissipated, optically thin molecular gas exists in regions with outflows (Dasyra et al., 2016; Oosterloo et al., 2017).
@@ 2 The Data
The level 1 data were reduced with version 1.12.5 of the JWST (Python-based) pipeline and the jwst_1202.pmap content of the Calibration Reference Data System (CRDS). In the first stage of the pipeline, the function ”Detector1” was run, which primarily corrects for detector artifacts and produces detection rates from time-dependent slopes. All parameters were set to their default values for exposures with two integrations. Then, the values of pixels with cosmic ray (CR) hits in one integration were replaced by their values in the other integration. For exposures with a single integration, ”Detector1” was run with the additional use of the ”find_showers” parameter, so that CRs could be identified. Their flux was modeled and subtracted from the relevant data by the pipeline. We performed two extra reduction steps of our own to improve the data products. The first step was the subtraction of the background from the science frames. A master background was created for each band (channel and sub-band) as the median of the four available exposures. It was then subtracted from each science frame of the same band. This two-dimensional background subtraction had the optimal result with respect to the removal of artifacts compared to the one-dimensional or three-dimensional subtraction that is offered at later stages of the pipeline. Our second additional step was the masking of bad pixels in the science data using a master bad pixel mask per band. We created this mask by merging the CRDS mask file with a mask of outliers (5) in the master sky frame. In the background-subtracted science data, the replacement values of bad pixels were computed by interpolation of their neighbors’ values using the ”Gaussian2DKernel” function of the ”astropy” package and a standard deviation of one pixel.
We input the background-subtracted data to the second stage of the pipeline, ”Spec2Pipeline”, for all relevant calibrations prior to the building of the 3-dimensional cube. These include the world coordinate system (wcs) information buildup, the flatfielding, the flux calibration, the straylight correction, the fringe removal, and the rectification of the data. Besides all functions that run by default, we also enabled the residual fringe removal function (and disabled the background subtraction function). During this step, we also created a version of the cubes without the straylight correction, to check for the impact of the cruciform artifact (see Section A.1).
@@
The integrated spectrum of all the emission in the MIRI mosaic of each band is shown in Fig. 2. An impressive number (39) of spectral lines is detected, most of which have spatially-resolved emission (see Section 3.1). Unlike the spectral lines, the mid-infrared continuum emission of IC5063 is nearly unresolved, originating in its largest fraction from the nuclear point-spread function (PSF; as seen in A.3). Due to flatfielding uncertainties in the long band of channel 4, its flux is on average lower (by 8%) than that indicated by the flux in previous channels. This calibration uncertainty also applies to fluxes of the lines in the same channel and band. It is comparable to the nominal flux calibration accuracy of the instrument (10%).
@@ 3.1 Overall emission properties and molecular disk properties
The continuum emission is largely unresolved, with more than 96% of its flux in all bands originating from the supermassive black hole vicinity, at 20:52:02.36-57:04:07.54, according to the location of the PSF in ch1. Its flux is 0.22(0.01), 0.95(0.02), 1.28(0.02), and 3.0(0.3) Jy at 6, 12, 14, and 25 respectively. In the MIRI wavelengths, we primarily see the AGN-heated dust, which is likely located in the black hole torus and which has been attributed a temperature of 720 K for its near infrared emission (Kulkarni et al., 1998). The torus has been predicted in the past from datasets of lower spatial resolution (e.g., Kulkarni et al. 1998; Hoenig et al. 2010; Esparza-Arredondo et al. 2019). It has been considered responsible for the detection of crepuscular rays in this galaxy (Maksym et al., 2020). Its maximum extent is limited, from the PSF in our ch1 data, to 50 pc, which improves the same measurement performed with ground-based mid-infrared sub-arcsecond observations (Asmus et al., 2014). A small contribution from the radio core to the nuclear spectrum is also possible, given that some continuum emission (up to 4%) exists between the nucleus and the radio lobes, which can be attributed to the jet-related synchrotron. Some weak silicate absorption is also seen at 9.7 , with optical depth =0.18, which indicates the existence of compact dust in or around the torus (and the radio core). The diffuse dust emission in the MIRI FOV is very weak compared to the nuclear point-source emission. Still, the spatially-resolved PAH emission, in particular that of the brightest 11.3 complex, confirms its presence.
@@
In contrast to the continuum emission, most of the line emission is spatially resolved. Thanks to the unprecedented quality of the MIRI data, 39 lines of molecular and atomic gas phase were detected, and spectral cubes were created for all of them except for the polycyclic aromaric hydrocarbons (PAHs). Images of the integrated line emission are shown in Fig. 3 and their total fluxes are presented in Table 1. The comparison with the previous Spitzer data is impressive. Less than half of the Table 1 spectral lines were previously present in spatially-integrated Spitzer Infra-Red Spectrograph (IRS) spectra (Tommasin et al., 2010; Panuzzo et al., 2011; Rampazzo et al., 2013), even though the IRS field of view was comparable to that of our MIRI mosaic for the short-high (SH) and long-high (LH) modes, and greater than that of our MIRI mosaic for the short-low (SL) and long-low (LL) modes. A bare minimum of kinematic information (i.e., a linewidth) was extracted for a handful of lines in the SH and LH modes (Guillard et al., 2012). Detailed kinematical analysis is now possible for nearly all lines, revealing out-of-dynamical-equilibrium components in many of them, e.g., in all of the previously detected or undetected lines - from S(1) to S(8). Depending on the line under examination, the emission probes the disk and regions within it or near it that are excited by the jet, or even the nucleus.
@@
The spectral line that best probes the disk extent and properties is S(1) (Figs. 3; 4). Faint spiral arms can be seen in all data from S(1) to S(3). In the upper part of the disk, a spiral arm starting north-east of the nucleus and heading north-west of it, can be seen in all upper panels of Fig. 4. In the lower part of the disk, a fainter spiral structure extending from south-west of the nucleus to south-east of it can be seen in Fig. 4 for molecular gas velocities in the range 30V200 km s-1 range. Spiral structure with similar orientation has been seen before in CO interferometric data (Dasyra et al., 2016, 2022). Given the detected disk rotation pattern that has been computed as the first moment of the S(1) cube (Fig. 4), then the upper side of the disk is closer to us if the spiral arms are trailing. The detected dust lane location, in the north part of the galaxy (e.g., Fig. 1), confirms this and verifies that the spiral arms are trailing. Several clumps bright in molecular gas, even brighter than those in the spiral arms, are seen along the radio jets axis: in the line connecting the north west (NW) to the south east (SE) radio lobe, the brightest clumps are seen near the radio lobes themselves for all lines, from S(1) to S(8). Still, other clumps also exist further away from the nucleus, in particular for the lower rotational number lines. For example, in the south-east, two bright discrete clumps are seen in S(1) 900 and 1200 pc away from the nucleus, for positive disk velocities (identified as clumps b1 and b2 in Fig. 5, where individual regions of interest are named).
@@
In contrast to the molecular gas, the ionized gas lines do not show any spiral structure at disk velocities. The emission of most these lines largely follows the jet direction from the nucleus all the way to the radio lobes, where intense jet cloud-interactions are known to be taking place (e.g., Oosterloo et al., 2000; Morganti et al., 2007; Tadhunter et al., 2014; Dasyra et al., 2015). The two radio lobes coincide with the bases of the two cones of the extended emission line region (EELR) (Morganti et al., 1998), which makes the galaxy appear roughly X-shaped in its large-scale ionized gas distribution in optical wavelengths (Colina et al., 1991). In the MIRI data, the brightest locations of the ionized gas emission are indeed the two the EELR bases (or radio lobes) and the nucleus. Lines of higher ionizational potential (IP) tend to have more nucleated emission than lines with low IP. For all lines with IPs 100 eV (shown after [Fe VII] in Fig. 3), more than two thirds of the emission is circumnuclear, linked to the AGN and to the radio core. From the radio lobes outwards, the emission splits up in two main branches (see, e.g., [Mg V] and [Fe VII] in Fig. 3 for the two branches north-west of the NW radio lobe and [S IV] and [Ne VI] for the branches south-east of the SE radio lobe), which is why we alternatively call the EELR bases as bifurcation points. Further out, these branches constitute the edges of the EELR, which is long known to be oriented at a position angle similar to that of the disk instead of perpendicular to it (Colina et al., 1991). Just like in the optical, the EELR orientation is similar to that of the jet in the mid-infrared. At distances greater than the bifurcation points, all of the ionized gas emission is constrained in the interior of the EELR. Good tracers of the jet-ionized gas along the jet trail and along the EELR edges are [Mg V] and [Ne VI], thanks to their high IPs (109.3 and 126.2 eV, respectively) and to the high resolution of MIRI at their wavelengths. We further explain the connection of the gas emission with the jet in the sub-section that follows.

### v2

@@ Abstract
We present James Webb Space Telescope MIRI data of the inner 32kpc2 of the galaxy IC5063, in which the jets of a supermassive black hole interact with the gaseous disk they are crossing. Jet-driven outflows were known to be initiated along or near the jet path and to modify the stability of molecular clouds, possibly altering their star formation properties. The MIRI data, of unprecedented resolution and sensitivity in the infrared, now reveal that there are more than ten discrete regions with outflows, nearly doubling the number of such known regions. Outflows exist near the radio lobes, at the nucleus, in a biconical structure perpendicular to the jet, and in a bubble moving against the disk. In some of them, velocities above escape velocity are observed. Stratification is also observed, with higher ionization or excitation gas attaining higher velocities. More outflows and bow shocks, found further away from the nucleus than the radio lobes, in regions without significant radio emission, reveal the existence of past or weak radio jets that interacted with the interstellar medium. The coincidence of the bow shocks with the optical extended emission line region (EELR) suggests that the jets also contributed to the gas ionization. Maps of the gas excitation temperature, Tex, indicate that the molecular gas is most excited in regions with radio emission. There, Tex is 100 K higher than in the EELR interior. We argue that a combination of jet-related shocks and cosmic rays is likely responsible for this excess molecular gas excitation.
@@ 1 Introduction
The advent of the James Webb Space Telescope (JWST), with its unprecedented spatial resolution and sensitivity in the infrared, enabled the astronomical community to make fast advancements in the field of galaxy evolution, by looking further in time for distant galaxies and deeper into nearby galaxies. Besides the detection of the first galaxies and even the first stars in them (e.g., Atek et al. 2022; Castellano et al. 2022; Naidu et al. 2022; Finkelstein et al. 2022, 2023; Donnan et al. 2023a, b; Harikane et al. 2023a; Leung et al. 2023; Maiolino et al. 2023a; Yan et al. 2023), a major galaxy evolution theme to be revised by the JWST is the impact of black holes (BHs) on galaxy evolution (including via outflows; Armus et al. 2022; Rupke et al. 2023; Cresci et al. 2023; Vayner et al. 2024; Leftley et al. 2024). Cycle 1 surveys already indicated that many more active galactic nuclei (AGN) exist early, above redshift 3-4 (Barro et al., 2023; Juodzbalis et al., 2023; Kocevski et al., 2023; Labbe et al., 2023; Lyu et al., 2023), compared to what was known from past observations in X-rays or other wavelengths, and that some black holes grew too fast too soon compared to the local BH mass vs. host galaxy relation (e.g., Harikane et al. 2023b; Larson et al. 2023; Maiolino et al. 2023b for 4z8.5).
@@
The turmoil that jets cause as they propagate through a dense ISM is demonstrated in this paper, with the aid of cycle 1 JWST data of the nearby, early-type galaxy IC5063 (taken for the program 2004; PI Dasyra). IC5063 is well suited for demonstrating the capabilities of the JWST, as it is a galaxy known for its numerous jet-cloud interactions. These interactions occur as radio jets cross a disk of dense gas for hundreds of parsecs, along a trajectory nearly parallel to the disk plane, prior to escaping it (Morganti et al., 1998). The jets collide with dense ISM clouds at the two radio lobes, 300-500 pc away from the nucleus, driving outflows. An outflow has been seen in absorption in HI in front of one of the radio lobes (Oosterloo et al., 2000; Morganti et al., 2007). Outflows have also been seen in emission in several regions, in tracers of all gas phases - atomic (Morganti et al., 2007; Dasyra et al., 2015; Congiu et al., 2017; Venturi et al., 2021) and molecular (Tadhunter et al., 2014; Morganti et al., 2015; Dasyra et al., 2015, 2016, 2022). In Very Large Telescope (VLT) SINFONI data, Dasyra et al. (2015) found numerous outflows starting at or near the jet trail: at least 6 discrete outflow starting points were reported for the central 22kpc2 of this galaxy, including at the two radio lobes. In some of these locations, gas moving faster than the escape velocity was detected. Moreover, the gas in nearly all of the inner 1.5 square kiloparsec of the galaxy shows distorted kinematic signatures due to the jet. In VLT MUSE data, Venturi et al. (2021) showed that, in the ionized gas as seen by [O III], the turbulence in the galaxy may reach 4 kpc in both directions perpendicular to the jet. Morganti et al. (2015) proposed even the propagation of a homonegeous jet-inflated cavity (a.k.a. cocoon) in the inner region of IC5063 based on the low rotational number J CO data distribution. The mass of all outflowing gas is at least 6106 , when taking into account the mass of the neutral gas and the cold molecular gas, using their most-abundant gas tracers, HI and CO (Morganti et al., 2007; Dasyra et al., 2016). In this computation, the cold gas mass was determined using a lower than Galactic CO luminosity to mass factor (Dasyra et al. 2016; see also Bolatto et al. 2013), because some of the molecular gas emission in the outflow’s location was found to be optically thin (Dasyra et al., 2016; Oosterloo et al., 2017), potentially related to cloud evaporation in the flow.
@@ 2 The Data
The level 1 data were downloaded with the ”astroquery” package of python, which queries the Mikulski Archive for Space Telescopes (MAST)111The pipeline-version of our data can be retrieved from MAST using the link https://doi.org/10.17909/jvw9-5x80 (catalog https://doi.org/10.17909/jvw9-5x80).. The data were reduced with version 1.12.5 of the JWST (Python-based) pipeline222This pipeline version can be retrieved from the link https://doi.org/10.5281/zenodo.10022973 . and the jwst_1202.pmap content of the Calibration Reference Data System (CRDS). In the first stage of the pipeline, the function ”Detector1” was run, which primarily corrects for detector artifacts and produces detection rates from time-dependent slopes. All parameters were set to their default values for exposures with two integrations. Then, the values of pixels with cosmic ray (CR) hits in one integration were replaced by their values in the other integration. For exposures with a single integration, ”Detector1” was run with the additional use of the ”find_showers” parameter, so that CRs could be identified. Their flux was modeled and subtracted from the relevant data by the pipeline. We performed two extra reduction steps of our own to improve the data products. The first step was the subtraction of the background from the science frames. A master background was created for each band (channel and sub-band) as the median of the four available exposures. It was then subtracted from each science frame of the same band. This two-dimensional background subtraction had the optimal result with respect to the removal of artifacts compared to the one-dimensional or three-dimensional subtraction that is offered at later stages of the pipeline. Our second additional step was the masking of bad pixels in the science data using a master bad pixel mask per band. We created this mask by merging the CRDS mask file with a mask of outliers (5) in the master sky frame. In the background-subtracted science data, the replacement values of bad pixels were computed by interpolation of their neighbors’ values using the ”Gaussian2DKernel” function of the ”astropy” package and a standard deviation of one pixel.
We input the background-subtracted data to the second stage of the pipeline, ”Spec2Pipeline”, for all relevant calibrations prior to the building of the 3-dimensional cube. These include the world coordinate system (wcs) information buildup, the flatfielding, the flux calibration, the straylight correction, the fringe removal, and the rectification of the data. Besides all functions that run by default, we also enabled the residual fringe removal function (and disabled the background subtraction function). During this step, we also created a version of the cubes without the straylight correction, to check for the impact of the cruciform artifact (see Section B.1).
@@
The integrated spectrum of all the emission in the MIRI mosaic of each band is shown in Fig. 2. An impressive number (39) of spectral lines is detected, most of which have spatially-resolved emission (see Section 3.1). Unlike the spectral lines, the mid-infrared continuum emission of IC5063 is nearly unresolved, originating in its largest fraction from the nuclear point-spread function (PSF; as seen in A.2). Due to flatfielding uncertainties in the long band of channel 4, its flux is on average lower (by 8%) than that indicated by the flux in previous channels. This calibration uncertainty also applies to fluxes of the lines in the same channel and band. It is comparable to the nominal flux calibration accuracy of the instrument (10%).
@@ 3.1 Overall emission properties and molecular disk properties
The continuum emission is largely unresolved, with more than 96% of its flux in all bands originating from the supermassive black hole vicinity, at 20:52:02.36 57:04:07.54, according to the location of the PSF in ch1. Given that target acquisition was performed on a nearby star (at 20:52:09.6647 57:04:15.41), the uncertainty in these coordinates is below 0.1: according to the JWST documentation333 https://jwst-docs.stsci.edu/, the telescope’s pointing accuracy with target acquisition is better than 0.1 and potentially as good as 0.014. The continuum flux is 0.22(0.01), 0.95(0.02), 1.28(0.02), and 3.0(0.3) Jy at 6, 12, 14, and 25 respectively. In the MIRI wavelengths, we primarily see the AGN-heated dust, which is likely located in the black hole torus. By fitting the nuclear emission at 2.2 and 3.8 , Kulkarni et al. (1998) deduced an indicative torus temperature of 720 K. By modeling the torus emission in lower spatial resolution data, Esparza-Arredondo et al. (2019) provided a torus mass range of 103-106 . The torus clumpiness and compactness have been considered responsible for the detection of crepuscular rays in this galaxy (Maksym et al., 2020). Its minimum extent, from predictions of the dust sublimation radius is 0.07 pc (Hoenig et al., 2010). Its maximum extent is limited, from the PSF our ch1 data, to 50 pc, moderately improving our knowledge from ground-based mid-infrared sub-arcsecond observations (82 pc; Asmus et al. 2014). A small contribution from the radio core to the nuclear spectrum is also possible, given that some continuum emission (up to 4%) exists between the nucleus and the radio lobes, which can be attributed to the jet-related synchrotron. Some weak silicate absorption is also seen at 9.7 , with optical depth =0.18, which indicates the existence of compact dust in or around the torus (and the radio core). The diffuse dust emission in the MIRI FOV is very weak compared to the nuclear point-source emission. Still, the spatially-resolved PAH emission, in particular that of the brightest 11.3 complex, confirms its presence.
@@
In contrast to the continuum emission, most of the line emission is spatially resolved. Thanks to the unprecedented quality of the MIRI data, 39 lines of molecular and atomic gas phase were detected, and spectral cubes were created for all of them except for the polycyclic aromaric hydrocarbons (PAHs). Images of the integrated emission of all lines are shown in the Appendix. Their total line fluxes are presented in Table 1. The comparison with the previous Spitzer data is impressive. Less than half of the Table 1 spectral lines were previously present in spatially-integrated Spitzer Infra-Red Spectrograph (IRS) spectra (Tommasin et al., 2010; Panuzzo et al., 2011; Rampazzo et al., 2013), even though the IRS field of view was comparable to that of our MIRI mosaic for the short-high (SH) and long-high (LH) modes, and greater than that of our MIRI mosaic for the short-low (SL) and long-low (LL) modes. A bare minimum of kinematic information (i.e., a linewidth) was extracted for a handful of lines in the SH and LH modes (Guillard et al., 2012). Detailed kinematical analysis is now possible for nearly all lines, revealing out-of-dynamical-equilibrium components in many of them, e.g., in all of the previously detected or undetected lines - from S(1) to S(8). Depending on the line under examination, the emission probes the disk and regions within it or near it that are excited by the jet, or even the nucleus. Images indicative of all features of interest are shown in Figs. 3, 5, and 4, for selected velocity ranges.
@@
The spectral line that best probes the disk extent and properties is S(1) (Fig. 3). Faint spiral arms can be seen in all data from S(1) to S(3). In the upper part of the disk, a spiral arm starting north-east of the nucleus and heading north-west of it, can be seen in all upper panels of Fig. 3. In the lower part of the disk, a fainter spiral structure extending from south-west of the nucleus to south-east of it can be seen in Fig. 3 for molecular gas velocities in the range 30V200 km s-1 range. Spiral structure with similar orientation has been seen before in CO interferometric data (Dasyra et al., 2016, 2022). Given the detected disk rotation pattern that has been computed as the first moment of the S(1) cube (Fig. 3), then the upper side of the disk is closer to us if the spiral arms are trailing. The detected dust lane location, in the north part of the galaxy (e.g., Fig. 1), confirms this and verifies that the spiral arms are trailing. Still, the brightest emission does not originate from the spiral arms, but from several clumps along the radio jet axis. In the line connecting the north west (NW) to the south east (SE) radio lobe, the brightest clumps are seen near the radio lobes themselves for all lines, from S(1) to S(8). Other bright clumps also exist further away from the nucleus, in particular for the lower rotational number lines. For example, in the south-east, two bright discrete clumps are seen in S(1) 900 and 1200 pc away from the nucleus, for positive disk velocities (identified as blobs b1 and b2 in Figs. 3 and 4).
@@
In contrast to the molecular gas, the ionized gas lines do not show any spiral structure at disk velocities (see the Appendix). The emission of most these lines largely follows the jet direction from the nucleus all the way to the radio lobes (see examples in Figs. 4 and 5, which separately present images for positive and negative velocity ranges so that regular and irregular kinematics can be easily distinguished). At the radio lobes, the bright emission is attributed to the intense jet cloud-interactions are known to be taking place (e.g., Oosterloo et al., 2000; Morganti et al., 2007; Tadhunter et al., 2014; Dasyra et al., 2015). The two radio lobes coincide with the bases of the two cones of the extended emission line region (EELR) (Morganti et al., 1998), which makes the galaxy appear roughly X-shaped in its large-scale ionized gas distribution in optical wavelengths (Colina et al., 1991). In the MIRI data, the brightest locations of the ionized gas emission are indeed the two the EELR bases (or radio lobes) and the nucleus. Lines of higher ionizational potential (IP) tend to have more nucleated emission than lines with low IP. For all lines with IPs 100 eV (shown after [Fe VII] in the first figure of the Appendix), more than two thirds of the emission is circumnuclear, linked to the AGN and to the radio core. From the radio lobes outwards, the emission splits up in two main branches (see, e.g., [Mg V] and [Fe VII] in the same figure, for the two branches north-west of the NW radio lobe and [S IV] and [Ne VI] for the branches south-east of the SE radio lobe), which is why we alternatively call the EELR bases as bifurcation points. Further out, these branches constitute the edges of the EELR, which is long known to be oriented at a position angle similar to that of the disk instead of perpendicular to it (Colina et al., 1991). Just like in the optical, the EELR orientation is similar to that of the jet in the mid-infrared. At distances greater than the bifurcation points, all of the ionized gas emission is constrained in the interior of the EELR. Good tracers of the jet-ionized gas along the jet trail and along the EELR edges are [Mg V] and [Ne VI], thanks to their high IPs (109.3 and 126.2 eV, respectively) and to the high resolution of MIRI at their wavelengths. We further explain the connection of the gas emission with the jet in the sub-section that follows.